# Benchmarking GPT-4 on Algorithmic Problems: A Systematic Evaluation of Prompting Strategies

### v1

@@ 3.   Reasoning Tasks
2 Operands
3 Operands
4 Operands
Nesting 2
((-21+47)*
@@ 4.1.   Prompting techniques
In the following, we present the prompting techniques involved in this study. Examples of each prompting technique on the three tasks are reported in paragraph 9.2 in the Appendix.
@@ 5.3.   Prompting methods and generalization
As previously noted, this general tendency is more evident on Arithmetic, where the performance gains deriving from prompting methods that produce explicit reasoning are greater overall. On the other hand, the performance gain from such prompting methods is more evenly spread across data splits in the case of ListOps, for which the GPT-4 performance on simple splits is already quite high even using Zero-shot prompting, as shown in Table 2 (see also paragraph 9.1 in the Appendix).
@@ 9.2.1.   ListOps
Prompt type
Example
Zero-shot
Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
A: The final result is (arabic numeral):
Zero-shot role
Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
A: The final result is (arabic numeral):
Few-shot
Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN37].
A: [MIN37]=3. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MAX[MIN41]2].
A: [MAX[MIN41]2]=2. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [SM[SM794][SM498]7].
A: [SM[SM794][SM498]7]=8. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
Symbolic CoT
Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [SM73].
A: [SM73]= 0. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [SM[SM86]1].
A: [SM[SM86]1]= [SM41]= 5. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX243]4[MAX937]].
A: [MIN[MAX243]4[MAX937]]= [MIN[MAX243]49]= [MIN449]= 4. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MAX[SM[SM85][MIN84]][SM3[MIN40]]].
Verbal CoT
Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN82].
@@
A: Let’s solve the following expression: [MIN82]. Simplifying the expression, we get to the final result: 2 Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[SM56][MAX87]].
A: Let’s solve the following expression: [MIN[SM56][MAX87]]. Simplifying an expression without nested parentheses, we get: [MIN[SM56]8]. Simplifying the expression, it becomes: [MIN18] Taking an immediate solution step, we get to the final result: 1. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MIN326]0[SM851]].
A: Let us recall the expression to be solved: [MIN[MIN326]0[SM851]]. By solving a simple expression, we obtain: [MIN[MIN326]04]. Solving a expression within a single pair of brackets, we get: [MIN204]. Taking an immediate solution step, we get to the final result: 0. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
Zero-shot CoT
@@
Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
A: Let’s think step-by-step.
@@ 9.2.2.   Arithmetic
Prompt type
Example
Zero-shot
@@
Q: Solve the following arithmetic expression computing the modulo 100 of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-66-(-84*(-34+0))).
A: The final result is (arabic numerals):
Zero-shot role
@@
Q: Solve the following arithmetic expression computing the modulo 100 of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-66-(-84*(-34+0))).
A: The final result is (arabic numerals):
Few-shot
@@
Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (51*39).
@@
A: (51*39)=89. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((28*-53)*(-76*90)).
A: ((28*-53)*(-76*90))=60. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (40-54-(-33–97+-19)).
A: (40-54-(-33–97+-19))=-59. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-66-(-84*(-34+0))).
Symbolic CoT
Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-16*-37).
A: (-16*-37)= 92. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((87*-51)-(47*-6)).
A: ((87*-51)-(47*-6))= ((87*-51)–82)= (-37–82)= 45. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-12–28-74)+-21+(76+-32+-87)).
A: ((-12–28-74)+-21+(76+-32+-87))= ((-12–28-74)+-21+-43)= (-58+-21+-43)= -22. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (((-33-39)*(67*65))-22).
Verbal CoT
Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-14*88).
A: Let us recall the expression to be solved: (-14*88). Simplifying the expression, we get to the final result: -32 Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((92*26)*(-35*59)).
@@
A: Let us recall the expression to be solved: ((92*26)*(-35*59)). Simplifying the expression, it becomes: ((92*26)*-65) Solving a expression within a single pair of brackets, we get: (92*-65). Simplifying the expression, we get to the final result: -80 Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (83-(46+-5-54)-25).
A: We need to solve the following expression: (83-(46+-5-54)-25). Taking an immediate solution step, we obtain: (83–13-25). As this expression is in a simple form, we can get to the final result: 71 Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-66-(-84*(-34+0))).
Zero-shot CoT
@@
Q: Solve the following arithmetic expression computing the modulo 100 of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (((-33-39)*(67*65))-22).
A: Let’s think step-by-step.
@@ 9.2.3.   Algebra
Prompt type
Example
Zero-shot
@@
Q: Simplify the following algebraic expression, computing the modulo 100 of the numerical coefficient of each intermediate value if it’s positive, and the modulo -100 if it’s negative: ((-78*b*x*y+(+50*b*x*y+-22*b*x*y))+((-b*x+-57*b*x)+(-38*b*x+-99*b*x))). If possible, factor by grouping the final result.
A: The final result is (algebraic expression):
Zero-shot role
@@
Q: Simplify the following algebraic expression, computing the modulo 100 of the numerical coefficient of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-85*a*y+((-98*a*x*y+2*a*x*y)+0*x*y*a)). If possible, factor by grouping the final result.
A: The final result is (algebraic expression):
Few-shot
@@
Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-55*b*x*y+-8*b*x). If possible, factor by grouping the final result.
@@
A: (-55*b*x*y+-8*b*x)=-b*x*(55*y+8). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-54*x*y+-68*x*y)+(-99*x*y++62*x*y)). If possible, factor by grouping the final result.
A: ((-54*x*y+-68*x*y)+(-99*x*y++62*x*y))=-59*x*y. Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((+12*x*y+-59*x*y++58*x*y)+(+36*x*y++13*x*y++93*x*y)+(+96*x*y+-55*x*y++73*x*y)). If possible, factor by grouping the final result.
A: ((+12*x*y+-59*x*y++58*x*y)+(+36*x*y++13*x*y++93*x*y)+(+96*x*y+-55*x*y++73*x*y))=+67*x*y. Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-85*a*y+((-98*a*x*y+2*a*x*y)+0*x*y*a)). If possible, factor by grouping the final result.
Symbolic CoT
Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+39*a*b*y++15*a*b*x*y). If possible, factor by grouping the final result.
A: (+39*a*b*y++15*a*b*x*y)= +3*a*b*y*(5*x+13). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+21*x*y+(-26*x*y+-92*x*y)). If possible, factor by grouping the final result.
A: (+21*x*y+(-26*x*y+-92*x*y))= (+21*x*y+-18*x*y)= +3*x*y. Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((+45*b*x++22*b*x+-47*b*x)+-62*b*x*y). If possible, factor by grouping the final result.
A: ((+45*b*x++22*b*x+-47*b*x)+-62*b*x*y)= (+20*b*x+-62*b*x*y)= -2*b*x*(31*y-10). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-78*b*x*y+(+50*b*x*y+-22*b*x*y))+((-b*x+-57*b*x)+(-38*b*x+-99*b*x))). If possible, factor by grouping the final result.
Verbal CoT
Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+10*a*b*x*y+-23*a*b*x). If possible, factor by grouping the final result.
A: The expression we need to solve is: (+10*a*b*x*y+-23*a*b*x). Simplifying the expression and factoring by grouping, we get to the final result: +a*b*x*(10*y-23) Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-8*a*x*y+(-38*a*x+-70*a*x)). If possible, factor by grouping the final result.
@@
A: Let us recall the expression to be solved: (-8*a*x*y+(-38*a*x+-70*a*x)). By solving a simple expression, we obtain: (-8*a*x*y+-8*a*x). As this expression is in a simple form, we can get to the final result factoring by grouping: -8*a*x*(y+1) Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+31*a*b*y+(-50*a*b*x+-64*a*b*x+-46*a*b*x)). If possible, factor by grouping the final result.
A: We need to solve the following expression: (+31*a*b*y+(-50*a*b*x+-64*a*b*x+-46*a*b*x)). Taking an immediate solution step, we obtain: (+31*a*b*y+-60*a*b*x). Taking an immediate solution step and factoring by grouping, we get to the final result: -a*b*(60*x-31*y). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-78*b*x*y+(+50*b*x*y+-22*b*x*y))+((-b*x+-57*b*x)+(-38*b*x+-99*b*x))). If possible, factor by grouping the final result.
Zero-shot CoT
@@
Q: Simplify the following algebraic expression, computing the modulo 100 of the numerical coefficient of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-85*a*y+((-98*a*x*y+2*a*x*y)+0*x*y*a)). If possible, factor by grouping the final result.
A: Let’s think step-by-step.

### v2

@@ 3.   Reasoning Tasks
2 Operands 3 Operands 4 Operands
Nesting 2 ((-21+47)*
@@ 4.1.   Prompting techniques
In the following, we present the prompting techniques involved in this study. Examples of each prompting technique on the three tasks are reported in paragraph 10.2 in the Appendix.
@@ 5.3.   Prompting methods and generalization
As previously noted, this general tendency is more evident on Arithmetic, where the performance gains deriving from prompting methods that produce explicit reasoning are greater overall. On the other hand, the performance gain from such prompting methods is more evenly spread across data splits in the case of ListOps, for which the GPT-4 performance on simple splits is already quite high even using Zero-shot prompting, as shown in Table 2 (see also paragraph 10.1 in the Appendix).
@@ 7.   Ethics statement
8.   Acknowledgements
We are grateful to OpenAI for granting free research access to the GPT-4 and GPT-3.5 APIs.
@@ 10.2.1.   ListOps
Prompt type Example
@@
Zero-shot Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]]. A: The final result is (arabic numeral):
@@
Zero-shot role Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]]. A: The final result is (arabic numeral):
Few-shot Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN37]. A: [MIN37]=3. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MAX[MIN41]2]. A: [MAX[MIN41]2]=2. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [SM[SM794][SM498]7]. A: [SM[SM794][SM498]7]=8. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
Symbolic CoT Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [SM73]. A: [SM73]= 0. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [SM[SM86]1]. A: [SM[SM86]1]= [SM41]= 5. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX243]4[MAX937]]. A: [MIN[MAX243]4[MAX937]]= [MIN[MAX243]49]= [MIN449]= 4. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MAX[SM[SM85][MIN84]][SM3[MIN40]]].
Verbal CoT Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN82]. A: Let’s solve the following expression: [MIN82]. Simplifying the expression, we get to the final result: 2 Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[SM56][MAX87]]. A: Let’s solve the following expression: [MIN[SM56][MAX87]]. Simplifying an expression without nested parentheses, we get: [MIN[SM56]8]. Simplifying the expression, it becomes: [MIN18] Taking an immediate solution step, we get to the final result: 1. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MIN326]0[SM851]]. A: Let us recall the expression to be solved: [MIN[MIN326]0[SM851]]. By solving a simple expression, we obtain: [MIN[MIN326]04]. Solving a expression within a single pair of brackets, we get: [MIN204]. Taking an immediate solution step, we get to the final result: 0. Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]].
Zero-shot CoT Q: MIN, MAX and SM are operators on lists of single-digit integers which have the semantics of minimum, maximum and sum modulo 10, respectively. Solve the following expression involving these operators: [MIN[MAX[MIN68]8][MAX[SM23]6]]. A: Let’s think step-by-step.
@@ 10.2.2.   Arithmetic
Prompt type Example
@@
Zero-shot Q: Solve the following arithmetic expression computing the modulo 100 of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-66-(-84*(-34+0))). A: The final result is (arabic numerals):
@@
Zero-shot role Q: Solve the following arithmetic expression computing the modulo 100 of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-66-(-84*(-34+0))). A: The final result is (arabic numerals):
@@
Few-shot Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (51*39). A: (51*39)=89. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((28*-53)*(-76*90)). A: ((28*-53)*(-76*90))=60. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (40-54-(-33–97+-19)). A: (40-54-(-33–97+-19))=-59. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-66-(-84*(-34+0))).
@@
Symbolic CoT Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-16*-37). A: (-16*-37)= 92. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((87*-51)-(47*-6)). A: ((87*-51)-(47*-6))= ((87*-51)–82)= (-37–82)= 45. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-12–28-74)+-21+(76+-32+-87)). A: ((-12–28-74)+-21+(76+-32+-87))= ((-12–28-74)+-21+-43)= (-58+-21+-43)= -22. Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (((-33-39)*(67*65))-22).
@@
Verbal CoT Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-14*88). A: Let us recall the expression to be solved: (-14*88). Simplifying the expression, we get to the final result: -32 Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((92*26)*(-35*59)). A: Let us recall the expression to be solved: ((92*26)*(-35*59)). Simplifying the expression, it becomes: ((92*26)*-65) Solving a expression within a single pair of brackets, we get: (92*-65). Simplifying the expression, we get to the final result: -80 Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (83-(46+-5-54)-25). A: We need to solve the following expression: (83-(46+-5-54)-25). Taking an immediate solution step, we obtain: (83–13-25). As this expression is in a simple form, we can get to the final result: 71 Q: Solve the following arithmetic expression taking each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-66-(-84*(-34+0))).
@@
Zero-shot CoT Q: Solve the following arithmetic expression computing the modulo 100 of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (((-33-39)*(67*65))-22). A: Let’s think step-by-step.
@@ 10.2.3.   Algebra
Prompt type Example
@@
Zero-shot Q: Simplify the following algebraic expression, computing the modulo 100 of the numerical coefficient of each intermediate value if it’s positive, and the modulo -100 if it’s negative: ((-78*b*x*y+(+50*b*x*y+-22*b*x*y))+((-b*x+-57*b*x)+(-38*b*x+-99*b*x))). If possible, factor by grouping the final result. A: The final result is (algebraic expression):
@@
Zero-shot role Q: Simplify the following algebraic expression, computing the modulo 100 of the numerical coefficient of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-85*a*y+((-98*a*x*y+2*a*x*y)+0*x*y*a)). If possible, factor by grouping the final result. A: The final result is (algebraic expression):
@@
Few-shot Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-55*b*x*y+-8*b*x). If possible, factor by grouping the final result. A: (-55*b*x*y+-8*b*x)=-b*x*(55*y+8). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-54*x*y+-68*x*y)+(-99*x*y++62*x*y)). If possible, factor by grouping the final result. A: ((-54*x*y+-68*x*y)+(-99*x*y++62*x*y))=-59*x*y. Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((+12*x*y+-59*x*y++58*x*y)+(+36*x*y++13*x*y++93*x*y)+(+96*x*y+-55*x*y++73*x*y)). If possible, factor by grouping the final result. A: ((+12*x*y+-59*x*y++58*x*y)+(+36*x*y++13*x*y++93*x*y)+(+96*x*y+-55*x*y++73*x*y))=+67*x*y. Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-85*a*y+((-98*a*x*y+2*a*x*y)+0*x*y*a)). If possible, factor by grouping the final result.
@@
Symbolic CoT Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+39*a*b*y++15*a*b*x*y). If possible, factor by grouping the final result. A: (+39*a*b*y++15*a*b*x*y)= +3*a*b*y*(5*x+13). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+21*x*y+(-26*x*y+-92*x*y)). If possible, factor by grouping the final result. A: (+21*x*y+(-26*x*y+-92*x*y))= (+21*x*y+-18*x*y)= +3*x*y. Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((+45*b*x++22*b*x+-47*b*x)+-62*b*x*y). If possible, factor by grouping the final result. A: ((+45*b*x++22*b*x+-47*b*x)+-62*b*x*y)= (+20*b*x+-62*b*x*y)= -2*b*x*(31*y-10). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-78*b*x*y+(+50*b*x*y+-22*b*x*y))+((-b*x+-57*b*x)+(-38*b*x+-99*b*x))). If possible, factor by grouping the final result.
@@
Verbal CoT Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+10*a*b*x*y+-23*a*b*x). If possible, factor by grouping the final result. A: The expression we need to solve is: (+10*a*b*x*y+-23*a*b*x). Simplifying the expression and factoring by grouping, we get to the final result: +a*b*x*(10*y-23) Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (-8*a*x*y+(-38*a*x+-70*a*x)). If possible, factor by grouping the final result. A: Let us recall the expression to be solved: (-8*a*x*y+(-38*a*x+-70*a*x)). By solving a simple expression, we obtain: (-8*a*x*y+-8*a*x). As this expression is in a simple form, we can get to the final result factoring by grouping: -8*a*x*(y+1) Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: (+31*a*b*y+(-50*a*b*x+-64*a*b*x+-46*a*b*x)). If possible, factor by grouping the final result. A: We need to solve the following expression: (+31*a*b*y+(-50*a*b*x+-64*a*b*x+-46*a*b*x)). Taking an immediate solution step, we obtain: (+31*a*b*y+-60*a*b*x). Taking an immediate solution step and factoring by grouping, we get to the final result: -a*b*(60*x-31*y). Q: Solve the following algebraic expression taking the numerical coefficient of each intermediate value modulo 100 if it’s positive, and modulo -100 if it’s negative: ((-78*b*x*y+(+50*b*x*y+-22*b*x*y))+((-b*x+-57*b*x)+(-38*b*x+-99*b*x))). If possible, factor by grouping the final result.
@@
Zero-shot CoT Q: Simplify the following algebraic expression, computing the modulo 100 of the numerical coefficient of each intermediate value if it’s positive, and the modulo -100 if it’s negative: (-85*a*y+((-98*a*x*y+2*a*x*y)+0*x*y*a)). If possible, factor by grouping the final result. A: Let’s think step-by-step.